OpenStudy (anonymous):

Solve by completing the square x^2+7x-8=0

OpenStudy (lgbasallote):

add 8 to both sides \[\implies x^2 + 7x = 8\] now divide 7 by 2 and square it \[\implies x^2 + 7x + \left ( \frac 72 \right) ^2 = 8 + \left(\frac 72 \right)^2\] simplify \[\implies x^2 + 7x + \frac{49}4 = 8 + \frac{49}4\] combine like terms... \[\implies x^2 + 7x + \frac {49}{4} = \frac{81}{4}\] factor out... \[\implies (x + \frac 72)^2 = \frac{81}4\] take square root of bot sides \[\implies x + \frac 72 = \pm \frac 92\] subtract both sides by 7/2 \[\implies x = \pm \frac 92 - \frac 72\] simpify to get final answer does that help?